### v1

@@ 4.2.2 Models allowing DM shape to vary from inner to outer regions
We create models with variable $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$ as specified in Eq. 5 to Eq. 8 for Auriga 12. We fix the transition radius $r_{q}=20$ kpc, while the remaining four parameters, $q_{\rm in,DM}$, $q_{\rm out,DM}$, $p_{\rm in,DM}$ and $p_{\rm out,DM}$, are treated as free parameters. The parameters determining the radial DM profile do not degenerate significantly with the 3D shape. To optimize the efficiency of the model, we fix them at $\log(\rho_{0}[M_{\odot}/{\rm kpc}^{3}])=6.2$, $r_{s}=40$ kpc, and $\gamma=1.6$.
The parameter grid we explored for this model is shown in Fig.10. Despite the significant uncertainties, a twisted DM halo being oblate ($p_{\rm in,DM}>q_{\rm in,DM}$ and $p_{\rm in,DM}\sim 1$) in the inner regions, and vertically aligned in the outer regions ($p_{\rm out,DM}<q_{\rm out,DM}$ and $q_{\rm out,DM}\sim 1$) is strongly favored.
We derive $p_{\rm DM}(r)$ and $p_{\rm DM}(r)$ from our models within the $1\sigma$ confidence level, and compare those with the true profiles derived from the simulation. Our model generally recovers the true tendency of the DM shape variation with radius, as shown in the right panels of Fig.10.
Such models allowing variable $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$ indeed are able to reproduce the stellar halo density distribution. We take one of the best-fitting models with $p_{\rm in}=1.04$, $p_{\rm out}=0.84$, $q_{\rm in}=0.6$, and $q_{\rm out}=1.32$, which reproduces the surface density distribution well from the inner to outer regions, with the residuals significantly reduced compared to the model with constant $p_{\rm DM}$ and $q_{\rm DM}$ as shown in Fig. 11. The tendency of $q_{\rm star}(r)$ derived from this model matches the data well (the dashed black curve in Fig. 9).
@@ 5 Discussion
The orientation of the Milky Way’s DM halo in the $x_{\rm gc}y_{\rm gc}$ plane remains largely undetermined. Ideally, if the DM orientation $\alpha_{q}$ could be constrained, we will converge at two solutions, one with $p<1$ and the correct $\alpha_{q}$ and one with $p^{\prime}=1/p$ when $\alpha_{q}$ is misaligned with 90. In our exploration of the Auriga 12 model allowing for a free orientation ($\alpha_{q}$) of the DM halo in the $x_{\rm gc}y_{\rm gc}$ plane, we failed to find significant constraints on $\alpha_{q}$. Notably, stellar density distributions created by various $\alpha_{q}$ values of the DM halo appeared similar, yielding no constraints on $\alpha_{q}$ from $\chi^{2}_{\rm den}$, and the same is true for the radial velocity distribution $v_{r}$. The velocities $v_{\phi}$ and $v_{\theta}$ are somewhat influenced by the DM orientation $\alpha_{q}$, but this effect occurs in a somewhat unpredictable manner.
This arises because we averaged out information along the azimuthal angle when correcting the selection function and comparing data with the model. We have to do so because observational data at various azimuthal angles remain incomplete, particularly with a lack of observations directed towards the Galactic Center. Consequently, this process results in a significant level of uncertainty for $p_{\rm DM}=Y/X$, especially when $p_{\rm DM}<q_{\rm DM}$, as demonstrated by Auriga 12. Despite this, the relative positioning of $p_{\rm DM}$ and $q_{\rm DM}$ is effectively constrained; in our models for Auriga 12, $p_{\rm DM}<q_{\rm DM}$ is strongly favored in the outer regions, in contrast to Auriga 23 and 5, where $p_{\rm DM}>q_{\rm DM}$ are strongly favored.
@@ 6 Summary
We present an empirical triaxial orbit-superposition model tailored for the Milky Way halo. In this method, we take the minimum assumption of dynamical models that the stellar halo is stationary, meaning the distribution function (DF) in 6D phase-space, $f(\boldsymbol{x},\boldsymbol{v})$, does not change when the stars are orbiting in the correct gravitational potential. The feasibility of this assumption can be gauged by the goodness of matching the data and the model at the end.
The gravitational potential of the model is a combination of disk, bulge, and DM halo with adaptable parameters. Our approach employs a highly flexible triaxial NFW halo, permitting variations of DM axis ratios ($p_{\rm DM}=Y/X$, $q_{\rm DM}=Z/X$) with radius. Within each specified gravitational potential, we take the 6D phase space information of stars observed as initial conditions, integrate their orbits in the gravitational potential, and form a model by superposing the orbits together. The orbits are weighted by the weights assigned to stars when correcting the observation selection function. We evaluate the goodness of the model by comparing the DF of the model and that directly constructed from the data, expressed in terms of density distribution and velocity distributions ($v_{r}$, $v_{\phi}$, $v_{\theta}$) across the $R_{\rm gc}-z_{\rm gc}$ plane.
This methodology is validated using mock observational data of the Milky Way halo from LAMOST + Gaia with three Auriga galaxies, halo 23, 5, 12 with different intrinsic shapes of DM halo. The mock data are created with similar spatial selection and observational errors as LAMOST + Gaia. We clean the data by subtracting the disk, substructures, etc., in exactly the same way as we deal with the real observational data.
@@
  • The density distribution and velocity distributions of the stellar halo provide consistent constraints on the DM distribution. Specifically, the stellar density distribution has stronger constraint on the DM shape, while velocity distributions better restrict the DM radial density profile.
  • The DM radial density profile and enclosed DM mass of the three galaxies within the data coverage are well recovered, with relative uncertainty of about $10\%$. Although we still have relatively large uncertainty in each of the three parameters determining the radial profile ($\rho_{0}$, $r_{s}$, and $\gamma$) with the limited data coverage.
@@
  • The DM halo of Auriga 12 is twisted being oblate ($p_{\rm DM}>q_{\rm DM}$) in the inner 20 kpc and vertically aligned ($p_{\rm DM}<q_{\rm DM}$) in the outer regions. Our model with constant $p_{\rm DM}$ and $q_{\rm DM}$ recovers those at the outer regions, and it does not reproduce the stellar halo density distribution in the inner regions well. We recover $p_{\rm DM}(r)$, $q_{\rm DM}(r)$ of Auriga 12 as a function of radius generally well with a model allowing them to vary, which also better reproduce the stellar halo density distribution from inner to outer regions.
@@ Acknowledgements.
We thank David Hogg, Jianhui Lian, Zhaozhou Li for useful discussions. LZ acknowledges the support from the CAS Project for Young Scientists in Basic Research under grant No. YSBR-062 and National Key R&D Program of China No. 2022YFF0503403. X.X.X. acknowledges the support from the National Key Research and Development Program of China No. 2024YFA1611902, CAS Project for Young Scientists in Basic Research grant No. YSBR-062 and YSBR-092, and the science research grants from the China Manned Space Project with NO. CMS-CSST-2025-A11.

### v2

@@ 4.2.2 Models allowing DM shape to vary from inner to outer regions
We create models with variables $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$ as specified in Eq. 5 to Eq. 8 for Auriga 12. We fix the transition radius $r_{q}=20$ kpc, while the remaining four parameters, $q_{\rm in,DM}$, $q_{\rm out,DM}$, $p_{\rm in,DM}$, and $p_{\rm out,DM}$, are treated as free parameters. The parameters determining the radial DM profile do not degenerate significantly with the 3D shape. To optimise the efficiency of the model, we fix them at $\log(\rho_{0}[M_{\odot}/{\rm kpc}^{3}])=6.2$, $r_{s}=40$ kpc, and $\gamma=1.6$.
The parameter grid we explored for this model is shown in Fig.10. Despite significant uncertainties, a twisted DM halo that is oblate ($p_{\rm in,DM}>q_{\rm in,DM}$ and $p_{\rm in,DM}\sim 1$) in the inner regions and vertically aligned in the outer regions ($p_{\rm out,DM}<q_{\rm out,DM}$ and $q_{\rm out,DM}\sim 1$) is strongly favoured.
We derive $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$ from our models within the $1\sigma$ confidence level, and compare those with the true profiles derived from the simulation. Our model generally recovers the true tendency of the DM shape variation with radius, as shown in the right panels of Fig.10.
Such models allowing variables $p_{\rm DM}(r)$ and $q_{\rm DM}(r)$ are indeed able to reproduce the stellar halo density distribution. We take one of the best fitting models with $p_{\rm in}=1.04$, $p_{\rm out}=0.84$, $q_{\rm in}=0.6$ and $q_{\rm out}=1.32$, which reproduces the surface density distribution well from the inner to outer regions, with the residuals significantly reduced compared to the model with constant $p_{\rm DM}$ and $q_{\rm DM}$ as shown in Fig. 11. The tendency of $q_{\rm star}(r)$ derived from this model matches the data well (the dashed black curve in Fig. 9).
@@ 5 Discussion
We have fixed the DM orientation with the tilt angle $\beta_{q}=0$. In reality, the DM halo might not be perfectly aligned with the stellar disk. The three galaxies we tested have DM tilt angles of $\beta_{q}\sim 10-15$. The axis ratio $q$ recovered from our model is close to the true axis ratio measured along the fixed $z_{\rm gc}$ axis, not the intrinsic $q$ measured along the principal axes of the DM halo, although the latter two are close to each other when $\beta_{q}$ is small. For galaxies with a larger tilt angle $\beta_{q}$, we expect the model to still recover $q$ measured along the fixed $z_{\rm gc}$ in a similar way; in this case, it may be significantly different from the intrinsic $q$ measured along the DM principal axes.
The orientation of the DM halo in the $x_{\rm gc}y_{\rm gc}$ plane remains largely undetermined. Ideally, if the DM orientation $\alpha_{q}$ could be constrained, we will converge to two solutions, one with $p<1$ and the correct $\alpha_{q}$ and one with $p^{\prime}=1/p$ when $\alpha_{q}$ is misaligned with 90. In our exploration of the Auriga 12 model that allows a free orientation ($\alpha_{q}$) of the DM halo in the $x_{\rm gc}y_{\rm gc}$ plane, we failed to find significant constraints on $\alpha_{q}$. In particular, the stellar density distributions created by various $\alpha_{q}$ values of the DM halo appeared similar, resulting in no constraints on $\alpha_{q}$ from $\chi^{2}_{\rm den}$, and the same is true for the radial velocity distribution $v_{r}$. The velocities $v_{\phi}$ and $v_{\theta}$ are somewhat influenced by the orientation of the DM $\alpha_{q}$, but this effect occurs in a somewhat unpredictable manner.
This arises because we averaged out information along the azimuthal angle when correcting the selection function and comparing the data with the model. We have to do so because observational data at various azimuthal angles remain incomplete, particularly with a lack of observations directed towards the Galactic Centre. Consequently, this process results in a significant level of uncertainty for $p_{\rm DM}=Y/X$, especially when $p_{\rm DM}<q_{\rm DM}$, as demonstrated by Auriga 12. Despite this, the relative positioning of $p_{\rm DM}$ and $q_{\rm DM}$ is effectively constrained. In our models for Auriga 12, $p_{\rm DM}<q_{\rm DM}$ is strongly preferred in the outer regions, in contrast to Auriga 23 and 5, where $p_{\rm DM}>q_{\rm DM}$ are strongly preferred.
Stellar halo of the three galaxies we tested is generally in dynamical equilibrium, but inevitably with some residual substructures and non-equilibrium features; they did not introduce obvious systematic bias according to our test results. The LMC, which fell in recently, may have perturbed the Milky Way halo and will affect the results of dynamical modelling. However, LMC mainly induces a bulk motion in the outskirts of the halo, while only mildly affects the inner halo, causing a velocity shift in $V_{z}$ with $\sim 10$ km/s at $r<50$ kpc (Erkal et al., 2020). Among the Auriga simulations, Auriga 25 has a recent minor merger with mass and orbits similar to LMC, but its perturbation to the disk and inner halo is significantly stronger than the LMC did to the Milky Way. We apply the method to Auriga 25 and show that even with some obvious non-equilibrium in the data, the DM distribution can still be roughly recovered but with larger statistical uncertainty (see the Appendix).
@@ 6 Summary
We present an empirical triaxial orbit-superposition model tailored for the Milky Way halo. In this method, we take the minimum assumption of dynamical models that the stellar halo is stationary, meaning that the distribution function (DF) in the 6D phase space, $f(\boldsymbol{x},\boldsymbol{v})$, does not change when the stars orbit in the correct gravitational potential. The feasibility of this assumption can be gauged by the goodness of the match of the data and the model in the end.
The gravitational potential of the model is a combination of disk, bulge, and DM halo with adaptable parameters. Our approach employs a highly flexible triaxial NFW halo, which allows variations in the DM axis ratios ($p_{\rm DM}=Y/X$, $q_{\rm DM}=Z/X$) with radius. Within each specified gravitational potential, we take the 6D phase-space information of stars observed as initial conditions, integrate their orbits in the gravitational potential, and form a model by superposing the orbits together. The orbits are weighted by the weights assigned to the stars when correcting the observation selection function. We evaluated the goodness of the model by comparing the DF of the model with that directly constructed from the data, expressed in terms of density distribution and velocity distributions ($v_{r}$, $v_{\phi}$, $v_{\theta}$) across the $R_{\rm gc}-z_{\rm gc}$ plane.
This methodology is validated using mock observational data of the Milky Way halo from LAMOST + Gaia with three Auriga galaxies, halo 23, 5, 12 with different intrinsic shapes of the DM halo. The mock data are created with similar spatial selection and observational errors as in LAMOST + Gaia. We clean the data by subtracting the disk, substructures, etc., in exactly the same way as we deal with the real observational data.
@@
  • The density distribution and velocity distributions of the stellar halo provide consistent constraints on the DM distribution. Specifically, the stellar density distribution has stronger constraints on the DM shape, while velocity distributions better restrict the DM radial density profile.
  • The DM radial density profile and the enclosed DM mass of the three galaxies within the data coverage are well recovered, with a relative uncertainty of about $10\%$. Although we still have relatively large uncertainty in each of the three parameters that determine the radial profile ($\rho_{0}$, $r_{s}$, and $\gamma$) with limited data coverage.
@@
  • The DM halo of Auriga 12 is twisted, oblate ($p_{\rm DM}>q_{\rm DM}$) in the inner 20 kpc, and vertically aligned ($p_{\rm DM}<q_{\rm DM}$) in the outer regions. Our model with constant $p_{\rm DM}$ and $q_{\rm DM}$ recovers those in the outer regions, and it does not reproduce the stellar halo density distribution in the inner regions well. We recover $p_{\rm DM}(r)$, $q_{\rm DM}(r)$ of Auriga 12 as a function of radius generally well with a model that allows them to vary, which also better reproduces the stellar halo density distribution from inner to outer regions.
@@ Acknowledgements.
We thank David Hogg, Jianhui Lian, Zhaozhou Li for useful discussions. LZ acknowledges the support from the CAS Project for Young Scientists in Basic Research under grant No. YSBR-062 and the National Key R&D Program of China No. 2022YFF0503403. X.X.X. acknowledges the support from the National Key Research and Development Program of China No. 2024YFA1611902, CAS Project for Young Scientists in Basic Research grant No. YSBR-062 and YSBR-092, and the science research grants from the China Manned Space Project with NO. CMS-CSST-2025-A11.
Appendix A The recovery of DM distributions for Auriga 25
Auriga 25 has a recent minor merger which significantly pertubated its stellar disk and inner halo. The infall orbit and mass of this merger are similar to those of the LMC in the Milky Way, but its pertubation to the disk and inner halo are significantly stronger than the LMC did to the Milky Way.
We orient Auriga 25 in the way that the LMC-like satellite is located in the southern hemisphere, and we take the particles in the north to create the mock data; the stars in the north look closer to dynamical equilibrium with less asymmetrical structures. The mock data are created in a similar way as for the other three galaxies. As shown in Figure 12, there are significant substructures in the inner halo. We removed the big tail at the end of the disk, but not the other substructures.
We applied the dynamical model to this galaxy in the same way as we did to the three galaxies we showed in the paper. In Figure 13, we show the recovery of the DM radial density distribution and the 3D shape of the DM halo. In this case, we have larger statistical errors in the results. However, the true DM distribution is still generally recovered by our model within the $1\sigma$ confidence level.
Figure 12: The surface brightness of Auriga 25 projected with different directions. Its disk was significantly destroyed by a recent merger. Both its stellar disk and its halo are not in fully dynamical equilibrium.
Figure 13: The recovery of DM radial density distribution and DM 3D shape of Auriga 25, similar to that shown in Figure 7 and Figure 8 for other galaxies.